an airplane has a speed of 120km/h in still air. the plane is headed 30 degrees east of north when a wind starts blowing due east at 30km/hr. What is the planes speed relative to the ground?
1
Expert's answer
2019-04-15T10:07:11-0400

Velocity of an airplane in still air:


"(120 \sin{30}, 120 \cos{30}) \dfrac{km}{h}"

Velocity of the wind is


"(30, 0) \dfrac{km}{h}"

Velocity of airplane relative to the ground is


"(120 \sin{30} +30, 120 \cos{30}) \dfrac{km}{h}"

 The plane's speed relative to the ground:


"v=\sqrt{(120 \sin{30} +30)^2+(120 \cos{30})^2} =137.5\dfrac{km}{h}"